OpenStudy (anonymous):

Solve x2 - 7x = -13. x equals quantity of negative 7 plus or minus I square root of 6 all over 2 x equals quantity of negative 7 plus or minus I square root of 3 all over 2 x equals quantity of 7 plus or minus I square root of 6 all over 2 x equals quantity of 7 plus or minus I square root of 3 all over 2

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@mathstudent55

mathslover (mathslover):

The given equation can be better written as : \(x^2 - 7x + \color{blue}{13} = -13 + \color{blue}{13} \\ x^2-7x+13 = \cancel{-13} + \cancel{\color{blue}{13}} \\ x^2 - 7x + 13 =0 \\ \) Now, use quadratic equation to find the roots of the simplified equation. \(\textbf{Roots} = \cfrac{-b \pm \sqrt{b^2-4ac}}{2a}\) Where the considered quadratic equation is : \(ax^2 + bx + c =0 \) Compare the above equation (\(ax^2 + bx+c =0\) ) with the simplified equation (\(x^2-7x+13=0\)) and find : a, b and c and put the values in the quadratic equation.

OpenStudy (anonymous):

-7 +- squareroot -7^2 - 4(1)(-13) / 2(1)

mathslover (mathslover):

I will write it what you wrote above : (for better understanding) \(\cfrac{-7 \pm \sqrt{(-7)^2 -4(1)(-13)}}{2(1)}\)

mathslover (mathslover):

Now, see, \(ax^2 + bx + c = 0\) \(x^2 \space - 7x + 13 = 0\) or \(x^2 + (-7x) + 13 =0 \) Compare the above two equations : a = 1 b = -7 c = 13

mathslover (mathslover):

You might have compared incorrectly. Rechec your process.

mathslover (mathslover):

*Recheck

OpenStudy (anonymous):

im really confused..

mathslover (mathslover):

Where are you confused @kiaracakes120513 ?

OpenStudy (anonymous):

i dont really know what i have to do @mathslover

mathslover (mathslover):

You did well to put a, b and c in the equation .. but, you predicted the value of "b" incorrectly. x^2 - 7x + 13 = 0 Here, b = - 7 not 7

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i did that

mathslover (mathslover):

Now put a, b and c again in the quad. equation

OpenStudy (anonymous):

\[-7 \pm \sqrt{-7^2 - 4(1)(-13)} \over 2(1)\] @mathslover

mathslover (mathslover):

See, it is -b in the numerator and we calculated b = - 7 so, -b becomes : -(-7) = 7

OpenStudy (anonymous):

and what happens after that? im still a little confused but im sorta getting it.. @mathslover

mathslover (mathslover):

It becomes : \(\cfrac{7 \pm \sqrt{49 - 52} }{2} \) => \(\cfrac{7 \pm \sqrt{-3}}{2}\)

mathslover (mathslover):

So, you get it as 7 plus minus square root of (-3) all over 2

OpenStudy (anonymous):

ohh i get it now

OpenStudy (anonymous):

can you help me with another please?

OpenStudy (anonymous):

5x2 = -30x - 65

mathslover (mathslover):

Transpose (-30x - 65) to LHS 5x^2 + 30x + 65 = 0 Take 5 common

OpenStudy (anonymous):

so do i put that in the quadratic formula?

mathslover (mathslover):

Yes! (After cancelling out from each term)

OpenStudy (anonymous):

\[-30\pm \sqrt{30^2-4(5)(65)} \over 2(5)\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

Yes, it is right Kiara.